Solve for both variables:
6x + 2y =10
5x -4y =-3
Teacher's directions: 1) solve each equation for same variable 2) set them equal to each other and solve for one variable 3) plug answer into both equations
4) write answer as (x,y)

Answer by amarjeeth123(569)

6x + 2y =10........equation 1
5x -4y =-3.........equation 2
From 1 we get 3x+y=5
y=5-3x
4y=20-12x.......equation 3
From 2 we get 4y=5x+3.......equation 4
Equating 3 and 4 we get,
20-12x=5x+3
12x+5x=20-3
17x=17
x=1
y=5-3=2
6(1)+2(2)=6+4=10
5(1)-4(3)=5-8=-3
(1,2) is the solution which satisfies both equations.